I need help graphing the slope. Its says graph the line described. Through (0,4); m= -1/2
I have to graph it but I don't know how thank you!

Answer by Alan3354(69443)

I need help graphing the slope. Its says graph the line described. Through (0,4); m= -1/2
-------------------
You don't graph a slope. You graph equations.
-----------
You can graph the equation that has a y-intercept of (0,4) and a slope of -1/2
----------
One way to do it:
Plot the y-int (0,4)
The graph's y value decreases 1 for each increase of 2 in the x direction.
That makes other points (2,3), (4,2), etc.
Plot at least 2 points, draw a line thru them.
-----------------
Another way:
Use the equation to find points.
(0,4); m= -1/2 --> y = (-1/2)x + 4
Pick values of x and find y.
Plot the points
Draw a line thru them